OpenStudy (anonymous):

Write the equation in point-slope form,slope intercept, and standard form, for the line that has a slope of 1/2& goes through the point (-5,9)

OpenStudy (anonymous):

\[(y1-y) = m(x1-x)\]

OpenStudy (anonymous):

Plug in your points for the y and x values and plug in 1/2 for the slope. Then solve as y=mx+b